### v1

# Biomedical Event Extraction via Structure-aware Generation

Abstract
Biomedical Event Extraction (BEE) is a critical task that involves modeling complex relationships between fine-grained entities in biomedical text data. However, most existing BEE models rely on classification methods that neglect the label semantics and argument dependency structure within the data. To address these limitations, we propose GenBEE, a generative model enhanced with a structure-aware prefix for biomedical event extraction. GenBEE constructs event prompts that leverage knowledge distilled from large language models (LLMs), thereby incorporating both label semantics and argument dependency relationships. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes with structural prompts, enriching the generation process with structural features. Extensive experiments on three benchmark datasets demonstrate the effectiveness of GenBEE and it achieves state-of-the-art performance on the MLEE and GE11 datasets. Furthermore, our analysis shows that the structural prefixes effectively bridge the gap between structural prompts and the representation space of generative models, enabling better integration of event structural information.
Index Terms:
Biomedical Event Extraction, Generative Models, Structural Information, Large Language Models, Knowledge Distillation
I Introduction
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks: (1) Biomedical Event Trigger Detection, which aims to identify trigger words and classify them by event types that represent the presence of biomedical events. (2) Biomedical Event Argument Extraction, which aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. Moreover, BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Moreover, many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for Event Extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem by utilizing encoder-decoder-based Pre-trained Language Models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works, such as GTEE [14] and AMPERE [15], have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
Figure 1: An illustration of the nested events identified from the given text. Nested events constitute approximately 25% of all events in biomedical text data.
However, most current generation-based EE models [14, 13, 15] have overlooked events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 provides an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). Moreover, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. The rest of the events are regarded as flat events. Identifying these complex structures is non-trivial, as it can reveal the underlying relationships between events. For example, a “Site2” argument in a “Binding” event is likely to be a trigger for a “Positive Regulation” event.
An intuitive way to incorporate structural information through prompt engineering [21] is by adding natural language prompts that describe complex structural information as part of the input to the generative PLM’s encoder. This helps elicit the rich knowledge distributed in the PLM. However, [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether generative PLMs, such as BART, contain sufficient domain knowledge to represent complex biomedical events. Therefore, how to effectively incorporate structural information into generative BEE models remains an open question.
To address the above challenges, we propose GenBEE (Generative Biomedical Event Extraction model with structure-aware prefix), a generative framework to fully exploit complex event structures in biomedical texts. GenBEE first constructs type description prompts to integrate the semantics of type labels. It then distills knowledge from large language models (LLMs), such as GPT-4 [23], to create event template prompts, incorporating additional weak supervision information, including dependencies among argument roles, to enhance contextual representation learning. Furthermore, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts, to infuse the generation process with structural features. Altogether, GenBEE serves as an end-to-end BEE framework that simultaneously tackles event trigger detection and event argument extraction, effectively leveraging the shared knowledge and dependencies between these two subtasks. The main contributions of this paper are summarized as follows:
  • We propose GenBEE, a structure-aware generation model for biomedical event extraction. Our end-to-end model can distill knowledge from the LLM to incorporate label semantics, dependency relationships among arguments, and shared knowledge between subtasks.
  • We introduce a structural prefix learning module to guide the generation process. This module constructs the structure-aware prefix to link continuous prompts with the representation space of generative models. Notably, this plug-and-play module can be added to any biomedical event extraction framework.
  • We evaluate GenBEE on three widely-used BEE benchmark datasets, namely MLEE, GE11, and PHEE. Extensive experimental results and analyses demonstrate the effectiveness of the proposed GenBEE model.
II Related Work
II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in the propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. However, these QA methods lack the dependency information between arguments, which could be helpful for assigning argument types.
Figure 2: The overall architecture of our proposed GenBEE model.
II-B Generative Event Extraction
Recently, another line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks involving augmented natural languages, used a variety of brackets and vertical bar symbols to represent embedded language labels, and tuned the model to predict these augmented natural language labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and do not consider incorporating semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model featuring label semantics with designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, providing more contextual information. However, not much work investigates the effectiveness of incorporating structural features of complex events into the generative models, even though this information has proven beneficial in some extractive models. Therefore, in this paper, we investigate methods to integrate complex event structural information into generative models.
III Method
In this section, we introduce our GenBEE model in detail. The proposed GenBEE model performs generative biomedical event extraction as follows: The extraction process is divided into a number of subtasks based on each pre-defined event type of a dataset. Given a Dataset $\mathcal{D}$ $=$ $\left\{\mathcal{C}_{j}\mid j\in[1,|\mathcal{D}|]\right\}$ with its pre-defined event type set $\mathcal{E}$ $=$ $\left\{e_{i}\mid i\in[1,|\mathcal{E}|]\right\}$, where $\mathcal{C}_{j}$ denotes a textual context and $e_{i}$ denotes an event type. Each subtask is denoted as $S_{e_{i},\mathcal{C}}$, for the type of event $e_{i}$ and the context $\mathcal{C}$, where $\mathcal{C}$ contains all the textual contexts in the dataset. As such, we have a total of $|\mathcal{E}|$ generation subtasks based on each event type for all textual contexts $\mathcal{C}$. The proposed model conducts training iteratively based on each subtask with all textual contexts in the dataset. Figure 2 shows the architecture of the proposed model, which consists of the following modules: Event Prompt Construction and Concatenation, Structural Prompt Construction, Prefix Generation, BART Encoding-Decoding, and Event Generation.
III-A Event Prompt Construction
TABLE I: Input event prompt examples for the GE11 dataset.
Event Type Type Description Event Template
Binding Binding events involve two or more molecules coming together to form a complex. Event trigger {Trigger} <SEP>  {Role_Theme} at binding site {Role_Site} and {Role_Theme2} at adjacent site {Role_Site2} form a complex, assisted by {Role_Theme3} and {Role_Theme4}.
Positive regulation Positive regulation events involve the activation of gene expression or signaling pathways. Event trigger {Trigger} <SEP>  Activator {ROLE_Cause} at control site {ROLE_CSite} initiates signaling at {ROLE_Site}, enhancing the function of {ROLE_Theme}.
Localization Localization events track the movement of a biological entity to a specific cellular or anatomical location. Event trigger {Trigger} <SEP>  From {ROLE_AtLoc}, {ROLE_Theme} relocates to {ROLE_ToLoc}.
Phosphorylation Phosphorylation events capture the enzymatic process of attaching a phosphate group to a target protein. Event trigger {Trigger} <SEP> Enzyme at {ROLE_Site} catalyzes the phosphorylation of {ROLE_Theme}.
This module constructs event prompts as part of the model input to the proposed GenBEE model for capturing the semantic meaning of event types and the dependency relationships between arguments. Event prompts not only provide semantic information to the model but also define the output format. The event prompt $\mathcal{N}_{e_{i}}$ for event type $e_{i}$ contains the following components:
  • Event Type - It specifies the expected event type to be extracted.
  • Type Description - It provides a description on the event type.
  • Event Template - It specifies the expected output format of the extracted event. The event template consists of three parts: a trigger, a separation marker <SEP> and a number of arguments. We use two types of placeholders, namely {Trigger} and {Role_<argument>}, to represent the corresponding trigger and arguments, respectively.
Table I shows some example event prompts for the GE11 dataset. For example, the event type description for the Binding event is “Binding involves two or more molecules coming together to form a complex for various biological processes and signaling pathways.”. Note that the type description is obtained according to the textual description of each event type provided in the technical report of the corresponding biomedical event extraction dataset (e.g., GE11 [20]). In the event template, the trigger part is “Event trigger {Trigger}”, which remains the same for all event types. The argument part is specific to each event type $e_{i}$. Each {Role_<argument>} serves as a placeholder for the corresponding argument role for the event. For example, the argument part for the Binding event in the GE11 dataset is: {Role_Theme} at binding site {Role_Site} and {Role_Theme2} at adjacent site {Role_Site2} form a complex, assisted by {Role_Theme3} and {Role_Theme4}.”.
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type from the biomedical event datasets, which requires domain knowledge. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the “Binding” event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles, constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
Figure 3: Prompts for eliciting knowledge from LLM to construct the event template for the “Binding” event in the GE11 dataset.
After constructing the event prompt $\mathcal{N}_{e_{i}}$ for the event type $e_{i}$, we concatenate it with the context $\mathcal{C}_{j}$. For each context $\mathcal{C}_{j}$, the input tokens $\mathcal{X}_{e_{i},j}$ in the proposed model is constructed as follows:
$\mathcal{X}_{e_{i},j}=[\mathcal{N}_{e_{i}};[SEP];\mathcal{C}_{j}]$ (1)
where “;” denotes the sequence concatenation, and $[SEP]$ is the special separator token in the PLM. For each subtask based on $e_{i}$, we construct the input tokens for all the contexts in the dataset.
III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
TABLE II: Example structural prompts for the Binding event.
Category Structural Prompts
General Events (S1) Explore events that frequently co-occur with <T>Binding<T> events, aiming to identify and analyze the interactions and dependencies among these events to enhance understanding of their interrelationships.
Overlapping Events (S2) Explore entities that serve as triggers in both <T>Binding<T> events and other event types, aiming to clarify the overlap in trigger roles across different contexts to better understand trigger versatility.
Nested Events (S3) Explore entities acting in multiple roles, including as roles in <T>Binding<T> events and differently in other events, highlighting the dynamics of role versatility and their implications for event structure.
(S4) Explore entities where the trigger of <T>Binding<T> events also acts as a role in other events, or vice versa, highlighting these complex inter-event relationships to identify patterns of event interaction.
After structural prompts are constructed, this module generates prefixes with an encoder-only PLM to embed structural information of biomedical events from the structural prompts. In particular, BioBERT is used to generate trainable soft prompts to the BART Encoding-Decoding module of GenBEE, guiding its generation process with event structural information.
First, structural prompts are concatenated with the separator token $[SEP]$ to construct a textual sequence $S_{e_{i}}$ as follows:
$S_{e_{i}}=\langle[CLS],S1_{e_{i}},[SEP],S2_{e_{i}},[SEP],...,S4_{e_{i}},[SEP]\rangle$ (2)
where $S1_{e_{i}}$ to $S4_{e_{i}}$ are structural prompts for event type $e_{i}$. Then, the pre-trained BioBERT is employed to encode the textual sequence $S_{e_{i}}$ for event type $e_{i}$ into dense vectors $Dense_{e_{i}}$ as follows:
$Dense_{e_{i}}=BioBERT({S}_{e_{i}})$ (3)
After obtaining the dense vectors $Dense_{e_{i}}$, we extract the representation $h_{[CLS]}$ from $Dense_{e_{i}}$, which is encoded from the [CLS] token. The [CLS] token is the start marker of the input to BioBERT. As the first vector in $Dense_{e_{i}}$, $h_{[CLS]}$ contains high-level event structural information distilled from BioBERT. We then input $h_{[CLS]}$ into a Feed Forward Neural Network (FFNN) to model the prefix $\mathcal{P}_{e_{i}}$ as follows:
$\mathcal{P}_{e_{i}}=FFNN(h_{[CLS]})$ (4)
The length $l$ of prefix $\mathcal{P}_{e_{i}}$ is a hyper-parameter, which is determined experimentally, to control the size of the soft prompts.
III-C Generative Model Training
The proposed GenBEE model employs a generative pre-trained language model BART [12] for encoding-decoding. In this module, the prefix $\mathcal{P}_{e_{i}}$ is integrated into both the encoder and decoder of the model. Specifically, from the prefix for event type $e_{i}$, we obtain the additional key and value matrices as $\mathcal{P}_{e_{i}}=K^{e_{i}}=V^{e_{i}}$, which can further be expressed as $K^{e_{i}}=\{k_{1}^{e_{i}},...,k_{l}^{e_{i}}\}$ and $V^{e_{i}}=\{v_{1}^{e_{i}},...,v_{l}^{e_{i}}\}$. Here, $k_{n}^{e_{i}}$ and $v_{n}^{e_{i}}$ $(n\in[1,l])$ are vectors with the same hidden dimension in the Transformer layer. These additional key and value matrices are concatenated with the original key and value matrices in the self-attention layers of the encoder and cross-attention layers of the decoder. As such, when computing the dot-product attention, the query matrix at each position is influenced by these structure-aware prefixes, which will subsequently influence the weightings of the representations for event generation. Note that the prefix $\mathcal{P}_{e_{i}}$ is learnable according to the querying event type $e_{i}$, thereby guiding the BART PLM to differentiate event types with different structural features.
Given the input tokens $\mathcal{X}_{e_{i},j}$ and the injected prefixes $\mathcal{P}_{e_{i}}$, the model first computes the hidden vector representation $H$ of the input $\mathcal{X}_{e_{i},j}$ with the injected prefixes $\mathcal{P}_{e_{i}}$ using a bidirectional BART Encoder:
$H=\text{Encoder}(\mathcal{P}_{e_{i}},\mathcal{X}_{e_{i},j})$ (5)
where each layer of the BART Encoder is a Transformer block with the self-attention mechanism.
The BART Decoder then takes in the hidden states $H$ from the BART Encoder to generate the text $\mathcal{Y}_{e_{i},j}=\{y_{1},\cdots,y_{n}\}$ ($n\in[1,|\mathcal{Y}_{e_{i},j}|]$) token by token (conditioned on its previous contexts). The injected prefix $\mathcal{P}_{e_{i}}$ and the hidden states $h^{D}$ of the decoder $D$ are also involved in the computation. More specifically, during the $n$-th step of generation, the autoregressive BART Decoder predicts the $n$-th token $y_{n}$ and computes the $n$-th hidden state $h_{n}^{D}$ of the decoder as follows:
$(y_{n},h_{n}^{D})=\text{Decoder}(y_{n-1},[\mathcal{P}_{e_{i}};H;h_{1}^{D},\ldots,h_{n-1}^{D}])$ (6)
Each layer of the BART Decoder is a Transformer block that includes two types of attention mechanisms: a self-attention mechanism and a cross-attention mechanism. The self-attention mechanism uses the decoder’s hidden state $h_{n}^{D}$ to reference the previously generated words in the decoder’s output sequence. The cross-attention mechanism uses both the encoder’s hidden state $H$ and the decoder’s hidden state $h_{n}^{D}$ to integrate contextual representation into the decoding process.
Then, this module outputs a structured sequence that starts with the start token “ <BOS>” and ends with the end token “ <EOS>”. Given the input sequence $\mathcal{X}_{e_{i},j}$ with the injected prefixes $\mathcal{P}_{e_{i}}$, the conditional probability $p(\mathcal{Y}_{e_{i},j}|\mathcal{P}_{e_{i}},\mathcal{X}_{e_{i},j})$ of the output sequence $\mathcal{Y}_{e_{i},j}$ is calculated progressively by the probability of each step as follows:
$p(\mathcal{Y}_{e_{i},j}|\mathcal{P}_{e_{i}},\mathcal{X}_{e_{i},j})=\prod_{n=1}^{|\mathcal{Y}_{e_{i},j}|}p(y_{n}|y_{1},\cdots,y_{n-1},\mathcal{P}_{e_{i}},\mathcal{X}_{e_{i},j})$ (7)
Finally, the BART Decoder calculates the conditional probability of its output and generates the sequence $\mathcal{Y}_{e_{i},j}$ for event type $e_{i}$ and context $\mathcal{C}_{j}$ as follows:
$\mathcal{Y}_{e_{i},j}=\text{Decoder}(\mathcal{P}_{e_{i}},\mathcal{X}_{e_{i},j})$ (8)
where $\mathcal{P}_{e_{i}}$ denotes the structural prefix for event type $e_{i}$ and $\mathcal{X}_{e_{i},j}$ denotes the input tokens for event type $e_{i}$ with context $\mathcal{C}_{j}$. After obtaining the output tokens $\mathcal{Y}_{e_{i},j}$, the proposed model can then perform training and inference.
For training, the trainable parameters of the GenBEE model include those in the encoder-only pre-trained language model BioBERT and the generative pre-trained language model BART. The training objective of the GenBEE model is to minimize the negative log-likelihood of the ground truth sequence $\mathcal{Y^{\prime}}_{e_{i},j}$, given the output sequence $\mathcal{Y}_{e_{i},j}$:
$Loss=-log\sum_{i=1}^{|\mathcal{E}|}P(\mathcal{Y^{\prime}}_{e_{i},j}|\mathcal{Y}_{e_{i},j})$ (9)
GenBEE is specifically trained to perform prediction based on the given event type. During training, it learns and incorporates the structural features of each event type. If the context $\mathcal{C}_{j}$ contains multiple events, the GenBEE model will generate output text for each event template, with each corresponding to a trigger and its associated argument roles. If the model does not give any prediction result on any triggers or argument roles for a given event type, the output will contain the corresponding placeholders only.
For inference, the proposed GenBEE model enumerates all event types and generates outputs for each event type based on the given context. After generating the output, the model compares the generated tokens with the specified event templates for each event type to identify the triggers and arguments accordingly. Finally, string matching [13] is employed to identify and capture the span offsets of the predicted triggers and arguments.
IV Experiments
TABLE III: Statistics of the datasets.
MLEE GE11 PHEE
Train Dev Test Train Dev Test Train Dev Test
# Documents 131 44 87 908 259 - - - -
# Sentences 1294 467 885 7926 2483 - 2897 965 965
# Events 3121 670 1894 10310 3250 - 3003 1011 1005
# Nested & Overlapping Events 773 397 315 2843 658 - 69 26 29
# Arguments 2887 1065 1887 6823 1533 - 15482 5123 5155
This section first describes the dataset, evaluation metrics, implementation details, and baseline models. Then, we present the experimental results to show the effectiveness of the proposed GenBEE model for biomedical event extraction.
IV-A Datasets
We have conducted experiments on the following three publicly available benchmark datasets, namely MLEE, GE11, and PHEE, for biomedical event extraction. MLEE [19] has 29 event types and 14 argument roles. We use the train/dev/test split given by the data provider. GE11 [20] has 9 event types and 10 role arguments. We use the train/dev/test split given by the shared task and evaluate the performance based on the development set, as the test set is unannotated and the official tool for evaluation is no longer available. PHEE [30] has 2 event types and 16 role arguments. We use the train/dev/test split provided by TextEE [31]. We preprocess each dataset by using the script of TextEE [31]. Table III presents the details of the splits and the datasets. For evaluation metrics, we follow the previous work [32, 10, 13] to adopt two metrics to evaluate performance. (1) Trigger Classification (Trg-C) F1-score: A trigger is correctly classified if the predicted span and the predicted event type match the gold ones. (2) Argument Classification (Arg-C) F1 score: An argument is correctly classified if the predicted span, event type, and role type match the gold ones.
IV-B Implementation Details
For a fair comparison with recent works, we leverage the BART-large [12] as the PLM in our proposed GenBEE model. For encoding structural prompts, we use a pre-trained BioBERT [33] model. We train GenBEE on an NVIDIA A100 40G GPU. The learning rate of BART-large is set to $10^{-5}$ and the learning rate of BioBERT is set to $10^{-6}$. Moreover, we train GenBEE for 50 epochs on PHEE and 80 epochs on MLEE and GE11. The batch size is set to 16 during training. The prefix length $l$ is set to 40, which is determined experimentally within the interval $l$ = {20, 30, 40, 50, 60}.
IV-C Baseline Models
We compare our GenBEE with the following EE models: (1) DyGIE++ [32] is a classification-based model that captures contextual information with the span graph propagation technique. (2) OneIE [34] is a classification-based model that utilizes global feature-aware graphs to capture cross-subtask and cross-instance interactions. (3) AMR-IE [35] is a classification-based model that captures syntactic characteristics in contexts using the AMR graph. (4) EEQA [36] is a classification-based model that formulates EE as a question-answering task. (5) TagPrime [37] is a state-of-the-art classification-based EE model that iteratively extracts event information based on each event type. (6) DEGREE [13] is a state-of-the-art generation-based EE model that integrates prompts for conditional generation. As the vanilla DEGREE model was designed mainly for general domain event extraction, we have re-implemented it using the prompts provided by [31] for the MLEE, GE11, and PHEE datasets. To ensure a fair comparison across models, we adopt the official codes of the above baselines and train them with the same data.
TABLE IV: Experimental results (%) for extracting events based on the MLEE, GE11 and PHEE datasets.
Methods Type PLM MLEE GE11 PHEE
Trg-C Arg-C Trg-C Arg-C Trg-C Arg-C
DyGIE++ (2019) Cls RoBERTa-l 80.6 65.8 67.1 62.8 70.1 53.9
OneIE (2020) Cls RoBERTa-l 80.9 65.2 67.3 63.9 69.8 52.0
EEQA (2020) Cls RoBERTa-l 79.3 65.3 66.6 62.7 70.3 53.1
AMR-IE (2021) Cls RoBERTa-l 80.2 66.4 67.9 63.3 69.7 53.5
TagPrime (2023) Cls RoBERTa-l 80.6 67.1 68.4 63.6 70.9 52.2
DEGREE (2022) Gen BART-l 78.0 64.6 65.2 60.5 67.6 51.4
GenBEE Gen BART-l 81.4 67.9 68.2 64.4 69.8 53.8
IV-D Experimental Results
Table IV shows the Trg-C F1-scores and Arg-C F1-scores based on the MLEE, GE11, and PHEE datasets. The best score is highlighted in bold, and the second-best score is underlined. Generation-based models and Classification-based models are indicated by “Gen” and “Cls”, respectively. The letter l in the column PLM denotes the large model. Overall, the GenBEE model achieves new state-of-the-art performance on the MLEE and GE11 datasets. Moreover, on the PHEE dataset, our model achieves the second-highest Arg-C F1-score among all baselines, while its Trg-C F1-score is still competitive with other classification-based models. Additionally, we reckon that the simpler event structures with less structural information in the PHEE datasets are the main reason that affects the performance of the GenBEE model.
Table IV also shows that GenBEE significantly outperforms the state-of-the-art generation-based model, DEGREE, in experiments based on the MLEE, GE11, and PHEE datasets. More specifically, GenBEE demonstrates an improvement of 3.4%, 3.0%, and 2.2% in Trg-C F1-score, and an improvement of 3.3%, 3.9% and 2.4% in Arg-C F1-score, over DEGREE on the MLEE, GE11, and PHEE datasets, respectively. We attribute GenBEE’s performance improvement to the use of event prompts and structural prefixes, which provide event structural information for biomedical event extraction.
For the experiments on the PHEE dataset, it is worth noting that the two state-of-the-art models, TagPrime and DyGIE++, are both sequence tagging models. We find that these two models mainly benefit from better span identification of trigger and argument words for achieving better performance. Even though the Trg-C F1-score of GenBEE slightly lags behind some baselines, GenBEE still achieves the second-highest Arg-C F1-score among all baselines. We attribute this to GenBEE’s end-to-end extraction style, which can be less susceptible to the error propagation issue that commonly affects pipeline methods.
TABLE V: Ablation study on the proposed GenBEE model based on different model configurations. We report numbers in F1-scores (%).
Model MLEE GE11 PHEE $\triangle$ Average
Trg-C Arg-C Trg-C Arg-C Trg-C Arg-C Trg-C Arg-C
GenBEE 81.4 67.9 68.2 64.4 69.8 53.8 - -
(1) -w/o Structural Prompts 80.8 67.2 67.5 63.5 69.6 53.4 -0.5 -0.7
(2) -w/o Structural Prefix 79.8 66.7 66.4 62.1 68.7 52.6 -1.5 -1.6
(3) -w/o Event Prompts 80.9 67.3 67.5 63.0 69.2 52.9 -0.6 -1.0
(4) -w/o (1) (2) (3) 78.0 64.6 65.2 60.5 67.6 51.4 -2.9 -3.2
V Analysis
V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different components on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following three components: structural prompts, structural prefixes, and event prompts. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that the structural prompts, structural prefixes, and event prompts significantly contribute to the performance of our proposed model, regardless of the dataset. Firstly, removing structural prompts results in a drop of 0.5% and 0.7% in the average Trg-C and Arg-C F1-scores, respectively. This demonstrates that structural prompts can boost the performance of the proposed model. Secondly, the removal of the structural prefix from GenBEE results in a reduction of 1.5% and 1.6% in the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this component in the model’s overall performance. Removing the event prompts further leads to a decrease of 0.6% and 1.0% in the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Overall, each component of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
V-B Performance Comparison between LLMs API and Fine-tuned PLMs
Given the immense potential of using large language models (LLMs) API with In-Context Learning (ICL) [21] across various NLP tasks under data efficient scenarios [38], we conduct experiments to compare the few-shot performance of the proposed GenBEE model with LLMs API. Specifically, we consider two widely used LLMs: GPT-4 [39] and Llama3-70b [40] for comparison. We access these LLMs through APIs from their official providers. As a part of the prompt, we provide LLMs with the same type of instructions, EE templates that we use in GenBEE, and a few ICL examples (positive ones). It is worth noting that the number of ICL examples is limited by the maximum context length supported by LLMs.
Table VI presents the performance results for LLMs and our proposed GenBEE model, categorized by shot numbers: 4-shot, 8-shot, 16-shot, and 32-shot. In the 4-shot learning scenario, the GPT-4 API achieves the highest F1 scores across all datasets, followed by the Llama3-70b API, with GenBEE showing the lowest performance. Similar trends are observed in the 8-shot and 16-shot learning scenarios, where all models exhibit performance improvements with increased learning data (shot numbers). However, for LLMs, adding more ICL examples beyond 16-shot does not necessarily result in further performance gains, as improvements tend to plateau or even slightly decline at higher shot levels. Notably, at 32-shot learning, GenBEE achieves the highest F1-scores across all datasets, particularly on the PHEE dataset with F1-scores of 43.7% and 38.5% for Trigger and Argument Classification, respectively. We infer that this is likely due to PHEE’s limited number of event types (only two), which reduces the number of unseen labels during few-shot training and allows GenBEE to leverage the provided examples more effectively. In comparison, datasets such as MLEE and GE11, with more diverse event types, present a greater challenge for the model, resulting in relatively lower performance gains. Generally, the fine-tuned GenBEE is able to outperform LLMs (using In-Context Learning) as the number of labeled data increases.
TABLE VI: Experimental results using LLMs API and GenBEE for few-shot learning based on the MLEE, GE11 and PHEE datasets. All reported numbers are in F1-scores (%).
Models MLEE GE11 PHEE
Trg-C Arg-C Trg-C Arg-C Trg-C Arg-C
4-shot
GPT-4 API 34.0 27.5 23.0 21.9 39.5 33.8
Llama3-70b API 31.8 26.7 19.3 20.5 36.5 30.5
GenBEE 11.1 8.4 5.3 5.5 21.0 14.5
8-shot
GPT-4 API 35.7 29.0 24.1 22.1 40.4 34.5
Llama3-70b API 33.0 28.5 20.3 20.7 37.4 31.8
GenBEE 18.0 14.2 9.0 7.9 29.8 24.6
16-shot
GPT-4 API 34.4 30.2 24.7 22.8 41.7 35.2
Llama3-70b API 33.2 30.0 20.6 22.0 37.8 32.0
GenBEE 27.6 22.8 18.4 14.3 38.6 33.7
32-shot
GPT-4 API 34.2 29.5 23.9 23.2 41.2 35.4
Llama3-70b API 32.9 29.8 20.5 21.8 36.9 31.5
GenBEE 35.7 31.4 25.2 24.3 43.7 38.5
Figure 4: A case study based on two examples taken from the GE11 dataset.
V-C Case Study
In this section, we compare GenBEE and DEGREE on two examples, which are presented in Figure 4, to illustrate the differences in their predictions. Example 1 presents a case where the DEGREE model incorrectly predicts the argument “Site2” in the Binding event triggered by “binding”, but our GenBEE model gives the correct prediction. We infer that our model has incorporated dependency information between arguments using event prompts for enhancing the performance of biomedical event extraction. Example 2 presents a case where the DEGREE model fails to predict the arguments “Theme” and “Site” in the Phosphorylation event, which is nested with the Binding event. However, the GenBEE model is able to predict the arguments of these two nested events correctly. We infer that our model has incorporated structural information using structural prefixes for effectively recognizing the relationships between nested events. As illustrated by the case study, with event prompts and structural prefixes, our proposed GenBEE model is able to perform biomedical event extraction effectively.
VI Conclusion
We propose GenBEE, a novel generative model with structure-aware prefixes for biomedical event extraction. The experimental results demonstrate that GenBEE outperforms strong baselines on the MLEE, GE11, and PHEE datasets. Moreover, our experiments show that prefixes effectively serve as a medium to link structural prompts with the representation space of generative models, thereby enhancing the model’s overall performance.